OpenStudy (anonymous):

Find the midpoint of side YZ,when Y(0,-4b) and Z (-2a,0)

jimthompson5910 (jim_thompson5910):

Average the corresponding coordinates Xmid = (x1 + x2)/2 = (0+(-2a))/2 = -2a/2 = -a Ymid = (y1+y2)/2 = (-4b+0)/2 = -4b/2 = -2b So the midpoint is (-a, -2b)